The perimeter of a pool tale is 30 ft. The table is twice as long as it is wide. What is the length of the pool table?
Answer by macston(5194)

.
W=width; L=length=2W
.
P=2(L+W)
30ft=2(2W+W)
15ft=3W
5ft=W The table is 5 feet wide
.
L=2W=2(5ft)=10ft
ANSWER: The table is 10 feet long.